OpenStudy (anonymous):

solve by elimination...2r-5s=9, 5r+2s=37

OpenStudy (anonymous):

2r-5s=9 multiply by 2 5r+2s=37 multiply by5 --------- 4r-10s=18 25r+10s=185 ------------ 29r=203 r=7

OpenStudy (anonymous):

I need an ordered pair for this